# Nested ensemble Kalman filter for static parameter inference in nonlinear state-space models

### v1

@@ 4.3 Nonlinear observation model
It has been assumed that the density, $f_{t}^{(\theta)}(\cdot|x)$, arising from the observation model is a linear and Gaussian conditional density; see (8). Suppose now that ${f}_{t}^{(\theta)}$ can be evaluated point-wise and may be neither linear nor Gaussian. Key to our approach in this scenario is that ${f}_{t}^{(\theta)}$ can be approximated by a linear Gaussian conditional density which we denote by $g_{t}^{(\theta)}(\cdot|x)$ and whose form is given by (8). For example, if ${f}_{t}^{(\theta)}(y|x)=\mathcal{N}(y_{t};h(x_{t}),R)$ for some nonlinear function $h(\cdot)$ (which may depend on $\theta$), an approximate density $g_{t}^{(\theta)}$ can be found be taking the first two terms of a Taylor series expansion of $h(\cdot)$ about the forecast mean $\mu_{t|t}$ (see e.g. evensen2022), or by augmenting $x_{t}$ to include $h(\cdot)$ in the forecast ensemble (see e.g. anderson01). An example of an appropriate $g_{t}^{(\theta)}$ in the case of a Gaussian observation model with state-dependent variance can be found in Section 5.2.
@@
Hence, the denominator $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$ can be computed analytically using Algorithm 3, for a suitable choice of $H$ and $R$, deduced from the example specific ${g}_{t}^{(\theta)}$. Parameter (and state) filtering may then proceed via the NEnKF (Algorithm 6) at the expense of inferential accuracy, due to the additional error induced by ${g}_{t}^{(\theta)}$.
Naturally, inferential accuracy can be maintained at the expense of computational cost, by eschewing the EnKF in favour of the particle filter with target ${p}_{t}^{(\theta)}(x_{t}|y_{0:t})$ given by (3). That is, Algorithm 1 can be run, for example, with a proposal density $q_{t}^{(\theta)}(x_{t}|x_{t-1},y_{t}):=p_{t}(x_{t}|x_{t-1})$. However, particles propagated myopically of the next observation can result in an observed-data likelihood estimator with high variance (delmoral15), which can be inefficient when used inside SMC2 (golikyp18). Methods which use the EnKF as a proposal mechanism inside a particle filter over states have been proposed by papadakis2010 (see also tamboli2021; leeuwen2019). In particular, use of $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ in place of $q_{t}^{(\theta)}(x_{t}|x_{t-1}^{a_{t-1}^{j}},y_{t})$ in Algorithm 1 corresponds to the weighted ensemble Kalman filter of papadakis2010. One possible implementation of this approach is to take $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ as the density associated with an ensemble member generated by the update step of Algorithm 3, for a suitable choice of $H$ and $R$, and conditional on the forecast ensemble $\tilde{x}_{t}^{1:N}$. The weight function then takes the form
@@
We therefore propose a third option, which leverages the inferential accuracy of SMC2 and the computational efficiency of the EnKF, and can be applied in scenarios where the latent state process is required on a finer time-scale than the observations. Here, the inner loop of SMC2 operates by using a particle filter over states, with target $\widetilde{p}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ given by (21) and a proposal mechanism $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ given by (22). Hence, in Algorithm 1, the propagate step samples $x_{t}^{j}\sim\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$, $j=1\ldots,N$, via a single run of Algorithm 3 with inputs $\theta$, resampled particles $x_{t-1}^{a_{t-1}^{1:N}}$ and a suitable choice of $H$ and $R$. It should be clear that the unnormalised weight of the $j$th particle $x_{t}^{j}$ is given by
@@
where $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$ is obtained from the single run of Algorithm 3 in the propagate step. We note that $p_{t}^{(\theta)}(x_{t}|x_{t-1})$ is sampled in step 1 of Algorithm 3 when generating the forecast ensemble, but need never be evaluated, since the corresponding forecast density, $\widehat{p}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$, appears in both $\widetilde{p}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ and $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$, and therefore cancels in the weight function. This necessarily sacrifices inferential accuracy (compared to targeting ${p}_{t}^{(\theta)}(x_{t}|y_{0:t})$) but allows for integrating $x_{t-1}$ (and $(x_{t,1},\ldots,x_{t,m})$ where relevant) out of the weight function. This procedure is known as Rao-Blackwellisation (RB) and can lead to weights with lower variance (see e.g. doucet2000). Finally, taking the average (over state particles $x_{t}^{1:N}$) of $\tilde{w}_{t}(\tilde{x}_{t}^{1:N},x_{t}^{j})$ in (4.3) estimates the observed-data likelihood $\widetilde{p}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$, that is, the denominator in (21). Such estimates can be used directly in the SMC2 scheme (see Algorithm 2) to give recursive exploration of a target which we denote by $\widetilde{\pi}_{\textrm{enkf},t}(\theta|y_{0:t})$. We refer to the resulting inference scheme for this target as RB-SMC2.
@@ 5 Applications
To illustrate the nested ensemble Kalman filter (NEnKF) and benchmark against existing methods, we consider four applications of increasing complexity. In Section 5.1, we perform inference on an Ornstein-Uhlenbeck model using synthetic data. We compare the performance of NEnKF against the augmented and particle ensemble Kalman filters (AEnKF and PEnKF of Sections 3.3 and 3.4). We additionally benchmark performance and accuracy using the output of an MCMC scheme which directly targets the parameter posterior. In Section 5.2, we apply the NEnKF and two implementations of SMC2 in the context of a nonlinear observation model, via a stochastic model of predator-prey interaction. Section 5.3 considers a real data example involving the spread of oak processionary moth in south-east England. Finally, in Section 5.4, we consider inference for the parameters governing a 10-dimensional Lorenz-96 model. For the second and final applications, inferential accuracy is benchmarked against the output of a particle MCMC scheme. All algorithms were coded in R and run on a desktop computer with an Intel Core i7-10700 processor and a 2.90GHz clock speed. For the Lorenz 94 application (Section 5.4), the per particle steps corresponding to reweighting, mutation and updating of $N$ were parallelised across 10 cores. The code is available from https://github.com/AndyGolightly/NestedEnKF.
@@ 5.1 Ornstein-Uhlenbeck process
We ran NEnKF with $M=1000$ particles and initial ensemble members $N=10$. A mutation step was triggered if the effective sample size (6) fell below 400 ($\gamma=0.4$). Updating of $N$ was triggered following a move step if the estimate $\hat{\sigma}_{N}^{2}$ of $\text{Var}[\widehat{p}_{\text{enkf},t}^{(\bar{\theta})}(y_{0:t}|\tilde{x}_{0:t}^{1:N})]$ exceeded 1.5, resulting in the update $\tilde{N}=\hat{\sigma}_{N}^{2}N$. We additionally ran NEnKF with a delayed acceptance kernel in the move step, based on the $k$-nearest neighbours surrogate of Section 4.2, with $k=3$. This gave a modest decrease in CPU time of around 10% over NEnKF; therefore all results are based on NEnKF with delayed acceptance.
We ran PEnKF and AEnKF with respective particle $M$ and ensemble member $N$ values chosen to give a CPU time in approximate agreement with that of NEnKF-DA. For PEnKF, we fixed $N$ at the average value obtained at termination of NEnKF over several replicate runs. Both schemes require choosing the hyper-parameters $a$ and $h$ used in the artificial evolution kernel. We followed liu2001 by setting $a=\sqrt{1-h^{2}}$ and $h=1-((3\delta-1)/2\delta)^{2}$ with the recommendation that the discount factor $\delta$ is around 0.950.990.95-0.99. We considered several choices of $\delta$ and report results for $\delta=0.97$, which we found to give best performance.
@@
Figure 2: OU example. Marginal parameter filtering means and 95% credible intervals under the NEnKF, PEnKF and AEnKF schemes. For reference, ground truth summaries obtained from the output of eMCMC are displayed at time 49.
@@
There are considerable differences in accuracy of AEnKF compared to NEnKF and PEnKF, with the former consistently over-estimating all posterior quantities of interest. This is particularly noticeable for the marginal posterior mean and standard deviation of $\log\theta_{3}$, which controls the intrinsic stochasticity of the latent process, and for which an assumed linear Gaussian relationship between this parameter and the observation process is likely to be unsatisfactory. There is comparatively little difference between NEnKF and PEnKF, although Figure 1 suggests that lower bias and RMSE is achieved for NEnKF.
@@ 5.2 Lotka-Volterra model
The Lotka-Volterra system admits a bivariate SDE giving a simple description of the time-course behaviour of prey ($X_{1,t}$) and predators ($X_{2,t}$). It has been ubiquitously used to benchmark competing inference algorithms (see e.g. Fuchs_2013; ryder21; graham22). The SDE takes the form
@@
where $H=(1,0)$. This (nonlinear) observation model effectively represents a Gaussian approximation to a Poisson model, and has been used in application in population ecology by lowe23 among others. Hence, in addition to vanilla SMC2 (see Section 2.3), we consider use of the EnKF as descsribed in Section 4.3, as appropriate for nonlinear observation models. In more detail, we implemented three schemes as follows:
@@
    Nested EnKF (NEnKF). The target at time $t$ is $\widehat{\pi}_{\text{enkf},t}(\theta|y_{0:t})$. Observed-data likelihood contributions $\widehat{p}_{\text{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$ are calculated using Algorithm 3 with $R$ replaced by $H\hat{\mu}_{t|t-1}$, permitting a linear Gaussian approximation of (27).
@@
    Rao-Blackwellised SMC2 (RB-SMC2). The target at time $t$ is $\widetilde{\pi}_{\text{enkf},t}(\theta|y_{0:t})$. Estimates of observed-data likelihood contributions $\widetilde{p}_{\text{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$ are calculated using Algorithm 1, with weight function (4.3) where ${f}_{t}^{(\theta)}(y_{t}|x_{t})$ is given by (27), $g_{t}^{(\theta)}(y_{t}|x_{t})=\mathcal{N}(y_{t},Hx_{t},R)$ and $\widehat{g}_{\text{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$ is obtained by running Algorithm 3, with $R=H\hat{\mu}_{t|t-1}$. To mitigate against a proposal distribution (for generating $x_{t}^{j}$ in Algorithm 1) that is potentially light-tailed compared to the target (21), we scaled $R=H\hat{\mu}_{t|t-1}$ by a factor of 2.
@@
Using the settings described above, 50 replicate runs of each scheme were performed, and the results summarised in Table 2 and Figure 3. We again compare the accuracy of each scheme by reporting bias and root mean square error (RMSE) of the estimators of the marginal posterior means and standard deviations of each $\log\theta_{i}$. For this application, we computed reference values of these quantities from a long run ($10^{6}$ iterations with $100$ state particles) of particle MCMC (see e.g. GoliWilk11, for an application of this scheme to the Lotka-Volterra SDE). Hence, for this application, we report accuracy with reference to the ‘ground truth’ marginal parameter posterior $\pi_{T}(\theta|y_{0:T})$, which is also the target of vanilla SMC2, albeit using a particle filter.

### v2

@@ 4.3 Nonlinear observation model
It has been assumed that the density, $f_{t}^{(\theta)}(\cdot|x)$, arising from the observation model is a linear and Gaussian conditional density; see (8). Suppose now that ${f}_{t}^{(\theta)}$ can be evaluated point-wise and may be neither linear nor Gaussian. Key to our approach in this scenario is that ${f}_{t}^{(\theta)}$ can be approximated by a linear Gaussian conditional density which we denote by $g_{t}^{(\theta)}(\cdot|x)$ and whose form is given by (8). That is, we take
[…]
for suitable choice of $\tilde{H}$ and $\tilde{R}$. For example, if ${f}_{t}^{(\theta)}(y|x)=\mathcal{N}(y_{t};h(x_{t}),R)$ for some nonlinear function $h(\cdot)$ (which may depend on $\theta$), an approximate density $g_{t}^{(\theta)}$ can be found be taking the first two terms of a Taylor series expansion of $h(\cdot)$ about the forecast mean $\mu_{t|t}$ (see e.g. Evensen et al., 2022). In this case, $\tilde{H}$ is a Jacobian matrix and we also include the offset term $h(\mu_{t|t})-\tilde{H}\mu_{t|t}$ in the observation mean. Alternatively, $x_{t}$ can be augmented to include $h(\cdot)$ in the forecast ensemble (see e.g. Anderson, 2001). Then, $x_{t}$ is replaced by $z_{t}=(x_{t}^{\prime},h(x_{t})^{\prime})^{\prime}$ and $\tilde{H}=(0_{d_{x}\times d_{x}},I)$ for an identity matrix $I$ of appropriate dimension. Examples of appropriate $g_{t}^{(\theta)}$ in the case of a Gaussian observation model with state-dependent variance can be found in Sections 5.2 and 5.3.
@@
Hence, the denominator $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$ can be computed analytically using Algorithm 3, with $H=\tilde{H}$ and $R=\tilde{R}$, deduced from the example specific ${g}_{t}^{(\theta)}$. Parameter (and state) filtering may then proceed via the NEnKF (Algorithm 6) at the expense of inferential accuracy, due to the additional error induced by ${g}_{t}^{(\theta)}$.
Naturally, inferential accuracy can be maintained at the expense of computational cost, by eschewing the EnKF in favour of the particle filter with target ${p}_{t}^{(\theta)}(x_{t}|y_{0:t})$ given by (3). That is, Algorithm 1 can be run, for example, with a proposal density $q_{t}^{(\theta)}(x_{t}|x_{t-1},y_{t}):=p_{t}(x_{t}|x_{t-1})$. However, particles propagated myopically of the next observation can result in an observed-data likelihood estimator with high variance (Del Moral and Murray, 2015), which can be inefficient when used inside SMC2 (Golightly and Kypraios, 2018). Methods which use the EnKF as a proposal mechanism inside a particle filter over states have been proposed by Papadakis et al. (2010) (see also Tamboli, 2021; van Leeuwen et al., 2019). In particular, use of $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ in place of $q_{t}^{(\theta)}(x_{t}|x_{t-1}^{a_{t-1}^{j}},y_{t})$ in Algorithm 1 corresponds to the weighted ensemble Kalman filter of Papadakis et al. (2010). One possible implementation of this approach is to take $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ as the density associated with an ensemble member generated by the update step of Algorithm 3, with $H=\tilde{H}$ and $R=\tilde{R}$, and conditional on the forecast ensemble $\tilde{x}_{t}^{1:N}$. The weight function then takes the form
@@
We therefore propose a third option, which leverages the inferential accuracy of SMC2 and the computational efficiency of the EnKF, and can be applied in scenarios where the latent state process is required on a finer time-scale than the observations. Here, the inner loop of SMC2 operates by using a particle filter over states, with target $\widetilde{p}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ given by (21) and a proposal mechanism $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ given by (22). Hence, in Algorithm 1, the propagate step samples $x_{t}^{j}\sim\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$, $j=1\ldots,N$, via a single run of Algorithm 3 with inputs $\theta$, resampled particles $x_{t-1}^{a_{t-1}^{1:N}}$, $H=\tilde{H}$ and $R=\tilde{R}$. It should be clear that the unnormalised weight of the $j$th particle $x_{t}^{j}$ is given by
@@
where $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$ is obtained from the single run of Algorithm 3 in the propagate step. We note that $p_{t}^{(\theta)}(x_{t}|x_{t-1})$ is sampled in step 1 of Algorithm 3 when generating the forecast ensemble, but need never be evaluated, since the corresponding forecast density, $\widehat{p}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$, appears in both $\widetilde{p}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ and $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$, and therefore cancels in the weight function. This necessarily sacrifices inferential accuracy (compared to targeting ${p}_{t}^{(\theta)}(x_{t}|y_{0:t})$) but allows for integrating $x_{t-1}$ (and $(x_{t,1},\ldots,x_{t,m})$ where relevant) out of the weight function. This procedure is known as Rao-Blackwellisation (RB) and can lead to weights with lower variance (see e.g. Doucet et al., 2000). Step $t$ of the Rao-Blackwellised particle filter is given by Algorithm 7.
Algorithm 7 Rao-Blackwellised particle filter (step $t$)
1. Resample. For $j=1,2,\ldots,N$, sample an index $a_{t-1}^{j}$ from $\{1,\ldots,N\}$ with probabilities $w_{t-1}^{1:N}$ and put $x_{t-1}^{j}:=x_{t-1}^{a_{t-1}^{j}}$.
2. Propagate. Run Algorithm 3 with inputs $\theta$, ensemble members $x_{t-1}^{1:N}$, $H=\tilde{H}$ and $R=\tilde{R}$ to obtain ${x}_{t}^{1:N}\sim\widehat{g}_{\textrm{enkf},t}^{(\theta)}(x_{t}|y_{0:t},\tilde{x}_{t}^{1:N})$ and $\widehat{g}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$.
3. Weight. For $j=1,2,\ldots,N$, construct and normalise the weight
$\tilde{w}_{t}(\tilde{x}_{t}^{1:N},x_{t}^{j})=\widehat{g}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})\times\frac{{f}_{t}^{(\theta)}(y_{t}|x_{t}^{j})}{{g}_{t}^{(\theta)}(y_{t}|x_{t}^{j})},\quad w_{t}^{j}=\tilde{w}_{t}(\tilde{x}_{t}^{1:N},x_{t}^{j})/\sum_{k=1}^{N}\tilde{w}_{t}(\tilde{x}_{t}^{1:N},x_{t}^{k}).$
 Output: likelihood estimate $\widehat{\widetilde{p}}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})=\frac{1}{N}\sum_{j=1}^{N}\tilde{w}_{t}(x_{t}^{1:N},x_{t}^{j})$ and weighted particles $\{x_{t}^{1:N},w_{t}^{1:N}\}$.
Finally, taking the average (over state particles $x_{t}^{1:N}$) of $\tilde{w}_{t}(\tilde{x}_{t}^{1:N},x_{t}^{j})$ in (4.3) estimates the observed-data likelihood $\widetilde{p}_{\textrm{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$, that is, the denominator in (21). Such estimates can be used directly in the SMC2 scheme (see Algorithm 2) to give recursive exploration of a target which we denote by $\widetilde{\pi}_{\textrm{enkf},t}(\theta|y_{0:t})$. We refer to the resulting inference scheme for this target as RB-SMC2. Given that the inner particle filter of RB-SMC2 corrects for the use of $g_{t}^{(\theta)}$ via reweighting, we anticipate better inferential accuracy than NEnKF (which uses $g_{t}^{(\theta)}$ without further correction), albeit at increased computational cost. Provided that a suitable $g_{t}^{(\theta)}$ can be sought, we expect that RB-SMC2 will provide greater computational efficiency than SMC2, albeit for a reduction in inferential accuracy. It therefore represents a compromise between NEnKF and SMC2. We investigate the performance of RB-SMC2 in Sections 5.2 and 5.3.
@@ 5 Applications
To illustrate the nested ensemble Kalman filter (NEnKF) and benchmark against existing methods, we consider four applications of increasing complexity. In Section 5.1, we perform inference on an Ornstein-Uhlenbeck model using synthetic data. We compare the performance of NEnKF against the augmented and particle ensemble Kalman filters (AEnKF and PEnKF of Sections 3.3 and 3.4). We additionally benchmark performance and accuracy using the output of an MCMC scheme which directly targets the parameter posterior. In Section 5.2, we apply the NEnKF and two implementations of SMC2 in the context of a nonlinear observation model, via a stochastic model of predator-prey interaction. Section 5.3 considers a real data example involving the spread of oak processionary moth in south-east England. Finally, in Section 5.4, we consider inference for the parameters governing a 10-dimensional Lorenz-96 model. For the second and final applications, inferential accuracy is benchmarked against the output of a particle MCMC scheme. All algorithms were coded in R and run on a desktop computer with an Intel Core i7-10700 processor and a 2.90GHz clock speed. For the Lorenz-96 application (Section 5.4), the per particle steps corresponding to reweighting, mutation and updating of $N$ were parallelised across 10 cores. The code is available from https://github.com/AndyGolightly/NestedEnKF.
@@ 5.1 Ornstein-Uhlenbeck process
We ran NEnKF with $M=1000$ particles and initial ensemble members $N=10$. A mutation step was triggered if the effective sample size (6) fell below 400 ($\gamma=0.4$). Updating of $N$ was triggered following a move step if the estimate $\hat{\sigma}_{N}^{2}$ of $\text{Var}[\log\widehat{p}_{\text{enkf},t}^{(\bar{\theta})}(y_{0:t}|\tilde{x}_{0:t}^{1:N})]$ exceeded 1.5, resulting in the update $\tilde{N}=\hat{\sigma}_{N}^{2}N$. We additionally ran NEnKF with a delayed acceptance kernel in the move step, based on the $k$-nearest neighbours surrogate of Section 4.2. We used $k=3$, which gave a good balance between computational efficiency and accuracy of the surrogate. This gave a modest decrease in CPU time of around 10% over NEnKF; therefore all results are based on NEnKF with delayed acceptance.
We ran PEnKF and AEnKF with respective particle $M$ and ensemble member $N$ values chosen to give a CPU time in approximate agreement with that of NEnKF-DA. For PEnKF, we fixed $N$ at the average value obtained at termination of NEnKF over several replicate runs. Both schemes require choosing the hyper-parameters $a$ and $h$ used in the artificial evolution kernel. We followed Liu and West (2001) by setting $a=\sqrt{1-h^{2}}$ and $h=1-((3\delta-1)/2\delta)^{2}$ with the recommendation that the discount factor $\delta$ is around 0.950.990.95-0.99. We considered several choices of $\delta$ and report results for $\delta=0.97$, which we found to give best performance.
@@
Figure 2: OU example. Marginal parameter filtering means and 95% credible intervals under the NEnKF, PEnKF and AEnKF schemes. For reference, ground truth summaries obtained from the output of MCMC are displayed at time 49.
@@
There are considerable differences in accuracy of AEnKF compared to NEnKF and PEnKF, with the former consistently over-estimating all posterior quantities of interest. This is particularly noticeable for the marginal posterior mean and standard deviation of $\log\theta_{3}$, which controls the intrinsic stochasticity of the latent process, and for which an assumed linear Gaussian relationship between this parameter and the observation process is likely to be unsatisfactory. There is comparatively little difference between NEnKF and PEnKF, although Figure 1 suggests that lower bias and RMSE is achieved for NEnKF. In the remaining applications, therefore, we focus on comparisons between NEnKF and SMC2 (and where appropriate, RB-SMC2).
@@ 5.2 Lotka-Volterra model
The Lotka-Volterra system admits a bivariate SDE giving a simple description of the time-course behaviour of prey ($X_{1,t}$) and predators ($X_{2,t}$). It has been ubiquitously used to benchmark competing inference algorithms (see e.g. Fuchs, 2013; Ryder et al., 2021; Graham et al., 2022). The SDE takes the form
@@
where $H=(1,0)$. This (nonlinear) observation model effectively represents a Gaussian approximation to a Poisson model, and has been used in application in population ecology by Lowe and Golightly (2023) among others. Hence, in addition to vanilla SMC2 (see Section 2.3), we consider use of the EnKF as descsribed in Section 4.3, as appropriate for nonlinear observation models. In more detail, we implemented three schemes as follows:
@@
    Nested EnKF (NEnKF). The target at time $t$ is $\widehat{\pi}_{\text{enkf},t}(\theta|y_{0:t})$. Observed-data likelihood contributions $\widehat{p}_{\text{enkf},t}^{(\theta)}(y_{t}|y_{0:t-1},\tilde{x}_{t}^{1:N})$ are calculated using Algorithm 3 with $R$ replaced by $\tilde{R}=H\hat{\mu}_{t|t-1}$, permitting a linear Gaussian approximation of (27).
@@
[…]
@@
Using the settings described above, 50 replicate runs of each scheme were performed, and the results summarised in Table 2 and Figure 3. We again compare the accuracy of each scheme by reporting bias and root mean square error (RMSE) of the estimators of the marginal posterior means and standard deviations of each $\log\theta_{i}$. For this application, we computed reference values of these quantities from a long run ($10^{6}$ iterations with $100$ state particles) of particle MCMC (see e.g. Golightly and Wilkinson, 2011, for an application of this scheme to the Lotka-Volterra SDE). Hence, for this application, we report accuracy with reference to the ‘ground truth’ marginal parameter posterior $\pi_{T}(\theta|y_{0:T})$, which is also the target of vanilla SMC2, albeit using a particle filter.